Professor Jones created a fictitious unit called the dulfinch (dulfin). If 1.75 dulfin = 1.5x10-2 m, then 88.8 dulfin2 is equivalent to _______ mm2.  


1
Expert's answer
2021-02-09T04:42:51-0500

1.75 dulfin = 1.5×10-2 m

1 dulfin = (1.5×10-2)/1.75 m

1dulfin2 = ("\frac{1.5×10^-²}{1.75}m" )2

1 dulfin2 = (0.85714×10-2m)2

1 dulfin2 = 0.734688×10-4 mm2

88.8 dulfin2 = 88×0.734688×10-4 mm2

88.8 dulfin2 = 65.24×10-4 mm2